Heather deposited $100 at the end of every month into an RRSP for 5 years. The      interest rate earned was 5% compounded daily. What was the accumulated value of the RRSP at the end of the 5 years? Round to the nearest cent.


1
Expert's answer
2021-09-13T16:31:39-0400

"d=\$100, r=5\%, k=365"

"PN=\frac{d((1+r/k)^{Nk}-1)}{r/k} =\frac{100((1+5/365 \times 1/100)^{365\times 5} -1)}{5/100\times 1/365 } =\$207322.05"